Consider the following problem: You are the chief
Answer #1

a)clock cycle timea

=(1/clock ratea)

=(1/2*109)

=0.5ns or 0.5 nanoseconds or 0.5*10-9s

c)clock cycle timeb

=(1/clock rateb)

=(1/8*109)

=0.125*10-9s or 0.125ns or 0.125nanoseconds

=0.5ns or 0.5*10-9s